A researcher wishes to conduct a study of the color preferences of new car buyers. Suppose that 60% of this population prefers the color red. If 12 buyers are randomly selected, what is the probability that between 33 and 77 (both inclusive) buyers would prefer red? Round your answer to four decimal places.
Answer #1

This is a binomial distribution question with
n = 12
p = 0.6
q = 1 - p = 0.4
where

x1 = 3 (I think you made a typo with your data)
x2 = 7
P(3< X <7) = P(X = 4) + P(X = 5) + P(X = 6)     (Assuming 3 and 7 on not inclusive

P(3< X <7) = 0.3195